OpenStudy (anonymous):

MEDAL AND FAN Which of the systems below is equivalent to the given system? 0.01 x - 0.05 y = 0.1 0.4 x + 0.3 y = 0.9 10 x - 5 y = 10 and 4 x + 3 y = 9 x - 5 y = 10 and 4 x + 3 y = 9 x - 5 y = 1 and 4 x + 3 y = 9

OpenStudy (xanthe):

T.T

OpenStudy (anonymous):

:)))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))))

OpenStudy (xanthe):

Which one looks the slightest bit closest to you?

OpenStudy (anonymous):

10 x - 5 y = 10 and 4 x + 3 y = 9

OpenStudy (xanthe):

Its trial and error multiply them by either 10,100 or 1000

OpenStudy (xanthe):

Since youve chosen the 1st option, my guess is youve multiplied the first equation by 1000 and the second by 10. So, it should look like 1000(0.01 x - 0.05 y = 0.1) = 10x - 50y=100 10(0.4 x + 0.3 y = 0.9) = 4x + 3y = 90

OpenStudy (anonymous):

yes

OpenStudy (xanthe):

But the answer doesnt resemble with that of the option given so, if you calculate them properly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x - 5 y = 10 and 4 x + 3 y = 9

OpenStudy (xanthe):

10x - 50y=100 is not the same as 10 x - 5 y = 10, so option 1 is incorrect

OpenStudy (xanthe):

Can you demonstrate your conclusion?

OpenStudy (anonymous):

im guessing because i need to do this rfast

OpenStudy (xanthe):

Ok, option 2 is the correct answer because if you were to multiply 100 and 10 to the respective equations; 100(0.01 x - 0.05 y = 0.1) 10 (0.4 x + 0.3 y = 0.9) you'll get the exact answer as the one you've given above x - 5 y = 10 and 4 x + 3 y = 9 .